A pressure transmitter has the following specifications:

Lower Range Limit = 0.5 MPa

Upper Range Limit = 10.5 MPa

Reference Accuracy = .05% of span

Span Error = .10% of span

Zero Error = .05% of span

Thermal Span Error = .10% of span

Overpressure Error = .09% of span

What is the total error budget in MPa? (Type in a four-place decimal).



1
Expert's answer
2021-03-23T08:04:16-0400

Find the total error budget as a percentage of the span:


"\epsilon=\pm(0.05\%+0.1\%+0.05\%+0.1\%+0.09\%)=\\=0.39\%"

The span is the difference between upper range limit (URL) and lower range limit (LRL):


"S=URL-LRL=10.5+0.5=10\text{ MPa}."


The total error budget in MPa is


"\epsilon_S=\epsilon·S=\frac{0.39\%·10}{100\%}=\pm0.039\text{ MPa}."